Mr Mukonda’s perfomance of his 170 Biostatistics class of students is given in an incomplete distribution below.


Variable. Frequence

0-10 12

10-20. 21

20-30 f1

30-40 40

40-50 f2

50-60 26

60-70 18

(a) If the median is 35. Find the missing frequences

(b) From Rusangu passing mark policy, explain if these results are normally distributed


1
Expert's answer
2021-01-26T03:24:55-0500

Solution:


"\sum f =170"

"12+21+f1+40+f2+26+18=170"

"117+f1+f2=170""f1+f2 =53 \space ...... (i)"


The median class is 30-40:


"M= L+ ({{N \over 2}-cf \over f}) i"

"35= 30 + ({85-(33+f1) \over 40})10""5={52-f1 \over 4}"

"20=52-f1 \therefore f1=32""f1+f2=53 \therefore f2=21"

"Ans: f1=32 \space and \space f2=21"


b). Since the data is clearly bell shaped with largest values lying in the middle while smallest values lie towards the extremes, we can assume that the data is normally distributed